### v7

# CT-AGRG: Automated Abnormality-Guided Report Generation from 3D Chest CT Volumes

CT-AGRG: Automated Abnormality-Guided Report Generation from 3D Chest CT Volumes
The rapid increase of Computed Tomography examinations have created a need for robust automated analysis techniques in clinical settings to assist radiologists managing their growing workload. Existing methods generate entire reports directly from 3D CT images, without explicitly focusing on observed abnormalities. This unguided approach can result in repetitive content or incomplete reports. We propose a new anomaly-guided report generation model, which first predicts abnormalities and then generates targeted descriptions for each. Evaluation on a public dataset demonstrates significant improvements in report quality and clinical relevance. We extend our work by conducting an ablation study to demonstrate its effectiveness.
1 Introduction
Three-dimensional Computed Tomography (3D CT) scans have emerged as indispensable tools in medical imaging [1]. Given the rapidly increasing number of scans to analyze [2, 3, 4], the time-consuming nature of the task [5] and the important demand for specialized radiological expertise in numerous healthcare systems [6, 7], computed-aided diagnosis and automating report generation have emerged as an active research area  [8, 9, 10]. Whether dealing with 2D [11, 12] or 3D [13, 14] images, numerous deep learning methods have been introduced to support healthcare professionals in tasks such as anomaly classification [13, 15], detection [16, 17, 18], segmentation [19, 20, 21], and report generation [9, 10, 22, 23]. Indeed, the availability of public datasets for 2D radiology images [12, 11] has facilitated the development of such methods [22, 23, 24, 25, 26, 27]. Recently, CT2Rep [10] has been presented as the first 3D method for automated report generation, trained and evaluated on the public CT-RATE dataset [14] of CT examination. CT2Rep, uses an Encoder-Decoder model to generate reports directly from 3D CT scans. However, this end-to-end approach may lead to incomplete reports [6]. Reading a CT study and composing a report [28], a radiologist detects abnormalities and then provides descriptions for each identified abnormality. Inspired by the radiologists’ workflow, we propose a method that first predicts anomalies and then generates a specific sentence for each anomaly. Our contributions are as follows:
  • We introduce a new abnormality-based sentence generation model that enhances chest 3D CT scan report generation performance, achievable with limited computational resources (single GPU, 24-hour training time).
  • We leverage a pre-trained language model with biomedical domain knowledge that we condition to generate a sentence for each abnormality.
  • We evaluate the model on a public dataset and we enhance our findings with an ablation study to show the effectiveness of each module.
2 Related work
2.1 2D Radiology report generation
Image captioning has emerged as a crucial task across diverse research domains [29, 30, 31, 32, 33]. This field has garnered significant attention in recent years in recent years [34, 35, 36], with applications in automated medical image reporting [37]. Early supervised methods for 2D radiology report generation [11, 12] primarily used Encoder-Decoder architectures [24, 25, 38] . The features of the input image are extracted using an encoder, either a CNN [39] or a Vision Transformer [40, 41]. These image-level features are then passed to a decoder, corresponding to a RNN [42] or Transformer Decoder [23, 26], which predicts the corresponding report. More recent methods introduce relational memory [23], pathology tags [43], or reinforcement learning [44] to improve the quality of generated reports. However, these methods generate the entire report only from the radiology image. Given the complexity and the inherent variability in radiological findings, these methods tend to produce incomplete reports [6]. Guided methods were then introduced to facilitate report generation [45, 46]. Descriptive sentences specific to identified regions or predicted abnormalities can be generated through the integration of an anatomical region detection or abnormality classification module [22].
Fig. 1: Overview of the method. Pre-training. The input volume $x$ is passed through a visual extractor $\Phi_{V}$ (either CT-ViT [47] or CT-Net [13]) to extract an embedding $h$. $h$ is then given to classification head $\Psi$ which predicts the logit vector $\hat{y}$. Step 1. $h$ is fed into 18 projection heads ($\Psi^{p}_{i},i\in\{1,\ldots,18\}$) followed by small classification heads ($\Psi^{c}_{i},i\in\{1,\ldots,18\}$), one for each label. This second step enables to obtain an embedding $h_{i}$ (and then $h^{a}_{i}$) specific to each label. Step 2. If a label indexed by $i$ is predicted as abnormal by its corresponding classification head $\Psi^{c}_{i}$, the associated embedding $h^{a}_{i}$ is transformed by a lightweight MLP $\Phi_{T}(e_{i})$ to obtain $e_{i}$. An abnormality-specific description is generated from a pre-trained GPT-2 using $e_{i}$.
2.2 3D Radiology report generation
New challenges arise with 3D imaging [48, 49]. Compared to 2D images, this modality of data introduces new difficulties due to the limited availability of public datasets [10] in this field and the significant computational resources needed [48, 49]. Abnormality detection  [18, 50] and classification [13, 14] for 3D CT scans have been the subject of research efforts in recent years. The Transformer-based visual feature extractor CT-ViT [47] has been introduced to deal with 3D CT scan efficiently, enabling the development of CT2Rep [10]. CT2Rep is the first method for automatic 3D CT scan radiology report generation, trained and evaluated on the CT-RATE public dataset [14]. CT2Rep is an auto-regressive model based on an encoder-decoder architecture [23], where visual features are extracted from the 3D CT scan using CT-ViT and then given to a Transformer Decoder [23, 41] that generates the report. To improve the completeness of generated reports, our approach builds upon previous work [22, 51] and introduces a step for abnormality classification to guide report generation based on previously predicted anomalies. This method offers greater control over the generation process.
3 Dataset
We use the CT-RATE public dataset [14] to train and evaluate our method. CT-RATE contains reconstructed non-contrast 3D chest CT volumes, corresponding reports and 18 distinct types of abnormalities extracted from reports using RadBERT [14, 52]. Train set has 17,799 unique patients (34,781 volumes), validation set has 1,314 patients (3,075 volumes) and test set has 1,314 patients (3,039 volumes). Following [13, 14], all volumes are either center-cropped or padded to achieve a resolution of $240\times 480\times 480$ with a in-slice spacing of $0.75$ mm and $1.5$ mm in the z-axis. Hounsfield Unit (HU) [53] values are clipped between $-1000$ to $+200$ before normalization to [-1, 1].
4 Method
As illustrated in Figure 1, we employ a two-stage approach for anomaly detection and description generation. Initially, we use a visual feature extractor pre-trained on a multi-label classification task. In the first stage, we perform multi-task learning with one classification head per anomaly. If an anomaly is detected, its associated vector representation is then passed to the second stage. Here, a pre-trained GPT-2 model generates a descriptive text of the identified anomaly.
4.1 Multi-label classification pre-training
The model receives a volume $x\in\mathbb{R}^{240\times 480\times 480}$ as input. This volume is passed to a visual feature extractor $\Phi_{V}$, either CT-Net [13] or CT-ViT [47]. CT-Net consists of ResNet [54] modules pre-trained on ImageNet [55] followed by a 3D convolutional network [56]. CT-ViT [47] is a Vision Transformer [40] based on the attention mechanism [41] computed from 3D patches extracted from the initial volume. From the initial volume $x$, both CT-Net and CT-ViT yield a vector representation $h\in\mathbb{R}^{2048}$, such that:
$h=\Phi_{V}(x).$ (1)
Then, $h$ is fed into a classification head $\Psi$ to predict $\hat{y}\in\mathbb{R}^{18}$, as follows:
$\hat{y}=\Psi(h)\,,$ (2)
where the components of the logit vector $\mathbf{\hat{y}}$ represent the anomaly scores for their respective class labels. The model is trained using a binary cross-entropy loss function [57, 58]. This initial step facilitates the training of a visual encoder $\Phi_{V}$ to extract salient visual features that enable accurate anomaly prediction.
4.2 Step 1: Multi-task single label classification
The multi-label classification head $\Psi$ of the pre-trained network is replaced by 18 independent projection heads [59], denoted as $\Psi^{p}_{i}$ ($i\in\{1,\ldots,18\}$), followed by classification heads denoted as $\Psi^{c}_{i}$ ($i\in\{1,\ldots,18\}$). The feature vector $h$ is first passed through each projection head $\Psi^{p}_{i}$ to obtain a unique embedding noted $h_{i}\in\mathbb{R}^{1024}$ per class label, such that:
$h_{i}=\Psi^{p}_{i}(h)\ \forall\ \ i\in\{1,\ldots,18\}\,.$ (3)
Subsequently, each embedding $h_{i}$ is fed into its corresponding classification head $\Psi^{c}_{i}$. The classification head outputs an anomaly score $\hat{y}_{i}\in\mathbb{R}$ and employs a binary cross-entropy loss $\mathcal{L}_{i}\in\mathbb{R}$ as the objective function. The prediction step can be formalized as:
$\hat{y}_{i}=\Psi^{c}_{i}(h_{i})\ \forall\ i\in\{1,\ldots,18\}.$ (4)
Here, $\hat{y}_{i}$ represents the predicted anomaly score for class $i$. The binary cross-entropy loss is then computed for each class, such that:
$\mathcal{L}_{i}(\hat{y}_{i},y_{i})=-[y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})],$ (5)
where $y_{i}\in\{0,1\}$ denotes the ground-truth label for class $i$.
For each projection $\Psi^{p}_{i}$ and classification $\Psi^{c}_{i}$ head, back-propagation [60] of the corresponding parameters is performed solely based on the associated loss $\mathcal{L}_{i}$. For the rest of the neural network, back-propagation is carried out using the sum of these losses, denoted as $\mathcal{L}\in\mathbb{R}$. Introducing this supervised multi-task learning step enhances classification performance [61] and enables the extraction of a unique vector representation $h_{i}$ for each abnormality $i$, conditioning sentence generation per abnormality in the subsequent step.
4.3 Step 2: Abnormality-based sentence generation
Abnormality selection A description is generated for each abnormality predicted by the classification module. During the training of this generation module, we consider only the labels correctly predicted as abnormal by the classification heads. During inference, we only consider labels predicted as abnormal by the corresponding classification heads. A label $i$ is classified as abnormal if its predicted score $\hat{y_{i}}$ exceeds a threshold computed from the validation set, which maximizes the F1-Score [62, 63, 64].
Abnormality feature extraction For each label indexed by $i$ and predicted as abnormal by its corresponding classification head $\Psi^{c}_{i}$, we extract and expend the corresponding embedding $h_{i}\in\mathbb{R}^{1024}$ to obtain a vector $h^{a}_{i}\in\mathbb{R}^{18\times 1024}$. This vector $h^{a}_{i}$ contains zeros for all components associated with other abnormalities, while retaining the original embedding $h_{i}$. Using a lightweight MLP noted $\Phi_{T}$, we project $h^{a}_{i}$ from the visual latent space to the textual latent space, yielding $e_{i}\in\mathbb{R}^{1024}$, such that:
$e_{i}=\Phi_{T}(h^{a}_{i})\,,$ (6)
where $\text{where }h^{a}_{i}=[\vec{0},\ \ldots,\ \vec{0},\ h_{i},\ \vec{0},\ \ldots,\ \vec{0}]\,,$
and 0 denotes a zero vector of dimension $\text{and }\vec{0}\text{ denotes a zero vector of dimension 1024.}$.
This allows for a focused representation of the predicted abnormality within the multi-abnormality embedding space, while maintaining the context of the original embedding.
Sentence generation For each label $i$ predicted as abnormal, we generate a descriptive sentence based on its associated embedding $e_{i}$. Inspired by prior work [22], the GPT-2 Medium Language model [65] fine-tuned on PubMed abstracts [66] is used for the Decoder part. GPT-2 is an auto-regressive model grounded in attention mechanism [41] where each embedded token only considers the contextual information provided by preceding tokens. Instead of using the conventional attention mechanism, the language model is conditioned on anomaly prediction by employing pseudo self-attention [67], noted $PS$. This approach injects the features associated with abnormality classification into the self-attention mechanism of GPT-2, such that:
$PS(e_{i},Y)=\text{softmax}\left((YW_{q})\begin{bmatrix}e_{i}w_{k}\\ YW_{k}\end{bmatrix}^{T}\right)\begin{bmatrix}e_{i}w_{v}\\ YW_{v}\end{bmatrix},$ (7)
where $w_{k}$ and $w_{v}$ are the projection parameters for key and value vectors [41] for $e_{i}$; $W_{k}$, $W_{q}$ and $W_{v}$ are the projection parameters for key, query and value for $Y$ which are textual embedded tokens. During training, the model is trained on a next token prediction task [68] using a cross-entropy loss function [10, 22, 23]. All parameters are frozen except for those of GPT-2. During inference, only the CT scan and the [BOS] token [65] are fed into the model, which generates a description for each predicted abnormality.
Report generation The final report is obtained by concatenating the generated sentences associated with each predicted anomaly.
NLG Metrics CE Metrics
Method METEOR ROUGEL BERT BART P R F1
CT2Rep [10] $0.173\text{\scriptsize$\pm 0.002$}$ $0.243\text{\scriptsize$\pm 0.002$}$ $0.865\text{\scriptsize$\pm 0.001$}$ $-3.314\text{\scriptsize$\pm 0.085$}$ $41.63\text{\scriptsize$\pm 2.01$}$ $38.12\text{\scriptsize$\pm 0.88$}$ $36.73\text{\scriptsize$\pm 0.76$}$
CT-AGRG w/ CT-ViT $0.190\text{\scriptsize$\pm 0.004$}$ $0.270\text{\scriptsize$\pm 0.002$}$ $0.864\text{\scriptsize$\pm 0.001$}$ $-2.936\text{\scriptsize$\pm 0.040$}$ $37.75\text{\scriptsize$\pm 0.46$}$ $55.35\text{\scriptsize$\pm 0.91$}$ $42.13\text{\scriptsize$\pm 0.23$}$
CT-AGRG w/ CT-Net $\mathbf{0.196}\text{\scriptsize$\pm 0.008$}$ $\mathbf{0.280}\text{\scriptsize$\pm 0.003$}$ $\mathbf{0.867}\text{\scriptsize$\pm 0.001$}$ $\mathbf{-2.714}\text{\scriptsize$\pm 0.179$}$ $\mathbf{45.74}\text{\scriptsize$\pm 0.35$}$ $\mathbf{62.59}\text{\scriptsize$\pm 0.14$}$ $\mathbf{50.11}\text{\scriptsize$\pm 0.15$}$
Ablation study
Multi-label
classification
Multi-task
classification
Multi-Abnormality
Embedding
Language
model
BLEU-4 METEOR R F1
$0.083\text{\scriptsize$\pm 0.003$}$ $0.127\text{\scriptsize$\pm 0.008$}$ $19.05\text{\scriptsize$\pm 0.19$}$ $23.80\text{\scriptsize$\pm 0.10$}$
$0.154\text{\scriptsize$\pm 0.008$}$ $0.186\text{\scriptsize$\pm 0.005$}$ $56.56\text{\scriptsize$\pm 0.12$}$ $46.11\text{\scriptsize$\pm 0.13$}$
$0.146\text{\scriptsize$\pm 0.005$}$ $0.189\text{\scriptsize$\pm 0.007$}$ $58.89\text{\scriptsize$\pm 0.12$}$ $47.22\text{\scriptsize$\pm 0.11$}$
$\mathbf{0.172}\text{\scriptsize$\pm 0.007$}$ $\mathbf{0.196}\text{\scriptsize$\pm 0.008$}$ $\mathbf{62.59}\text{\scriptsize$\pm 0.14$}$ $\mathbf{50.11}\text{\scriptsize$\pm 0.15$}$
Table 1: Quantitative evaluation on the test set with Natural Language Generation (NLG) metrics and Clinical Efficacy (CE) metrics. Reported mean and standard deviation metrics were computed over five independent runs. Our model outperforms the previous method both on NLG and CE metrics. Ablation study: The ablation study on CT-AGRG was conducted with CT-Net [13] as visual feature extractor, since it yields the best results. It highlights the importance of each module.
5 Experimental setup
Training details. For the first and second steps, the model was trained with a batch size of 4, using the Adam [69] optimizer for 40 and 2 epochs, with a learning rate of $10^{-4}$ and $10^{-5}$, respectively. For the third step, training was conducted for 50 epochs with a batch size of 64, using the AdamW [70] optimizer with a learning rate of $5\times 10^{-5}$. The training required a GPU with 48GB of memory.
Generation mode. Generated sentences were limited to 60 tokens [71], typically the maximum observed per sentence in the report corpus. At inference, we use the Beam Search [72, 73, 74] algorithm as generation mode [75, 76, 77, 78]. Beam size is set to 4, aligning with prior work and balancing output quality with computational efficiency. [10, 22, 23].
6 Results
6.1 Quantitative results
Abnormalities mentioned in the generated report are extracted with the RadBERT labeler [14, 52]. These predicted anomalies are compared against ground truth labels to compute Clinical Efficacy (CE) metrics: Precision, Recall, and F1-Score [79]. We also evaluate our captioning model using widely used [10, 22, 23, 26] Natural Language Generation (NLG) metrics: BLEU-4 [80] for n-gram precision; METEOR [81] for alignment with stemming and synonymy; ROUGE [82] for overlap focusing on recall; BERT-Score [83] for token-level contextual embedding similarity; and BART-Score [84] for sequence-level likelihood assessment. This combination captures lexical, semantic, and generative aspects of caption quality, providing a comprehensive evaluation of our model’s performance [85, 86]. Table 1 demonstrates that CT-AGRG achieves significantly better results than prior work. With CT-ViT and CT-Net as visual feature extractors, we achieve an average Recall of 55.35 and 62.59 ($\Delta$+45.20and $\Delta$+64.19increase compared to CT2Rep [10]), and an average F1-Score of 42.13 and 50.11, respectively ($\Delta$+14.70and $\Delta$+36.43increase). CT-AGRG achieves a BART-Score [84] of -2.714, outperforming CT2Rep [10] (-3.314, $\Delta$+18.11increase), indicating enhanced abnormality detection capabilities and improved semantic relevance in generated captions.
6.2 Ablation study
As shown in Table 1, we conduct an ablation study to evaluate each component of our CT-AGRG model. Introducing multi-task classification significantly improves F1-Score from 23.80 to 46.11 ($\Delta$+93.76increase). Augmenting the latent space further enhances performance, increasing Recall from 56.56 to 58.89 ($\Delta$+4.12increase) and F1-Score from 46.11 to 47.22 ($\Delta$+2.41increase). NLG metrics show notable improvements, with BLEU-4 [80] and METEOR [81] scores increasing by 0.089 and 0.069, respectively, indicating enhanced sentence variability while maintaining semantic fidelity to ground truth reports.
7 Conclusion
In this paper, we have presented an effective method for guided report generation. Our approach, based on conditional description generation for each abnormality, outperforms state-of-the-art methods while offering a high level of control over the generated report. Future work could focus on adapting the approach to require less supervision, leveraging other modalities or region-specific data to further enhance the report generation process.
8 Acknowledgments
We acknowledge Hamamci et al. 2024 [14] for making the CT-RATE dataset available.
Appendix A Appendix: Qualitative results
Fig. 2: Comparison of ground truth with reports generated by CT2Rep and CT-AGRG from the CT-RATE test set.
Fig. 3: Comparison of ground truth with reports generated by CT2Rep and CT-AGRG from the CT-RATE test set.